OpenStudy (anonymous):

\[\frac{ d }{ dx }\left[ (x ^{2} +1)\ln x\right]\]

hartnn (hartnn):

do you know product rule of derivatives ?

OpenStudy (raden):

use the multiply rule : y ' = u' v + uv'

OpenStudy (anonymous):

yes

OpenStudy (raden):

oppss. product not multiply.. haha :)

OpenStudy (anonymous):

so wouldn't it be

OpenStudy (anonymous):

\[2x \times \ln x +(x ^{2}+1)\frac{ 1 }{ x }\]

hartnn (hartnn):

yes, thats correct ! :)

OpenStudy (anonymous):

oh nevermind lol the book just simplified it even more into \[(x ^{2}+1)/x+2xlnx\]

hartnn (hartnn):

okk..but its the same thing! :O

OpenStudy (anonymous):

yeah the way they wrote it was weird i couldn't figure out why 2xlnx was on the bottom and then i realized it wasn't they are just adding it ot the other one that is a fraction